OpenStudy (awkwardpanda):

Given A= [4, 9, -2, -3, 5, 7] and C= [2, -1, -3, 1, 4, -2], and A+B=C, what is B? a. B=[6, 8, -5, -2, 9, 5] b. B=[2, 10, 1, -4, 1, 9] c. B=[-2, -10, -1, 4, -1, -9] d. B=[-6, -8, 5, 2, -9, -5]

OpenStudy (awkwardpanda):

2. IF A=[3, -2, 1, 4] and B=[0. -1, 3, 2], what is AB? a. [0, 2, 3, 8] b. [3, -3, 4,6] c. [-1, -4, 11, 2] d. [-6, -7, 12, 7]

OpenStudy (awkwardpanda):

3. given a matrix A=[2, 3, -5, -7] what is A^-1? a. no inverse b.[-7, -5, 3, 2] c. [-7, -3, 5, 2] d. [7, 3, -5, -2]

OpenStudy (anonymous):

For the first one, B = C - A; so, just subtract the components of A from C.

OpenStudy (awkwardpanda):

so it would be B?

OpenStudy (anonymous):

I don't think so. From the first component of C and A, we get -2.

OpenStudy (awkwardpanda):

oh it's negative, whoops xD number 2?

OpenStudy (anonymous):

Number 2 seems odd to me because in order to multiply two matrices, the columns of the first matrix must be equal to the rows of the second. Moreover, if we are calculating the product of two vectors, their dot product will be a number.

OpenStudy (anonymous):

The only suggestion I could make would be A.

OpenStudy (awkwardpanda):

idk why i thought it was C ._. number 3?

OpenStudy (anonymous):

If you don't mind me asking, what course is this for?

OpenStudy (awkwardpanda):

Algebra 2.

OpenStudy (anonymous):

For number 3, there is no inverse since A is not n x n

OpenStudy (awkwardpanda):

YES I GOT THAT RIGHT. Thank you :D would you mind helping with others?

OpenStudy (anonymous):

Sure. I must go soon, so I will do as many as I can.

OpenStudy (awkwardpanda):

alright. \[\left\{ \frac{ -32x-15y+2z=1001 }{ 73x-3y+811z=-29 } \right\}\] there's 3 rows. the 3rd row is -51x+32y-513z=416

OpenStudy (anonymous):

Is this a system of linear equations?

OpenStudy (awkwardpanda):

i believe so..

OpenStudy (anonymous):

OK. So I find it easiest to use Gauss-Jordan Elimination to reduce the system to reduced row echelon form. Do you know how to do this?

OpenStudy (anonymous):

These numbers are unnecessarily large.

OpenStudy (awkwardpanda):

I don't remember xD i took this 3 months ago.. they are i agree.

OpenStudy (anonymous):

Another question, are you sure the problem was written down correctly?

OpenStudy (awkwardpanda):

Yes. There are 3 rows but the equation thingy only let me write 2

OpenStudy (anonymous):

Ok. Well, to go through the process of this particular problem would be long and painful because of the numbers involved. The solution is here, http://www.wolframalpha.com/input/?i=-51x%2B32y-513z%3D416+%3B+%E2%88%9232x%E2%88%9215y%2B2z%3D1001%3B+73x%E2%88%923y%2B811z%3D%E2%88%9229 .

OpenStudy (anonymous):

Check out this page for how to approach these problems. I must go now. Good luck in your studies. http://www.purplemath.com/modules/systlin6.htm

OpenStudy (awkwardpanda):

Thank you.